OpenStudy (anonymous):

Please Help!!! select the factored form of the expression. if it can not be factored, select "already in factored form." 6x^2-36 A) (6x-6)(6x+6) B) 6(x^2-6) C) already in factored form D)6x(x-6)

OpenStudy (anonymous):

Let's factor it! Take out a 6, so we have 6(x^2 - 6). Can we take a step further? We can, but that would result in square-roots. x^2 - 6 is actually = (x - √6)(x + √6), but that's not a choice. So the answer would have to be B.

OpenStudy (anonymous):

You could also just look at the answer choices, and eliminate the ones that don't make sense. It can't be A. (6x-6)(6x+6) because when you're going to multiply them, you'll get 36x^2, and that's not in our expression. It can't be D. either because 6x(x-6) = 6x^2 - 36x, which isn't our expression. We're left with B and C.

OpenStudy (anonymous):

oh ok i get it now thanks!! @TurtleMuffin